A missle luncher fire a missle at an angle of 43 above the horizontal has muzzle velocity of 200m/s a tank is advancing towards the luncher on a level ground at a speed of 10 m/s what should the distance from the luncher to the tank at the at the instant the luncher is firedin order to hit the tank g(10).


1
Expert's answer
2021-02-28T07:39:51-0500
"t=\frac{2v\sin{43}}{g}=2\frac{200\sin{43}}{9.8}=27.8\ s"

"s=v\cos{43}t+ut\\s=(200)\cos{43}(27.8)+10(27.8)=4344\ m"